# SEED: Customize Large Language Models with Sample-Efficient Adaptation for Code Generation

### v2

@@ III-C Model Optimization
For fine-tuning [24], we update the entire parameter $\theta$ of LLMs as
$\theta^{*}=\arg\min_{\theta}\sum_{(r,c^{*})}l_{ce}(\mathcal{M}_{\theta}(r),c^{*}),$ (6)
When the computational resources are insufficient, we employ Low-Rank Adaptation (LoRA) [25] to fine-tune LLMs. For a weight matrix $W\in\mathbb{R}^{d\times k}$, LoRA represents its update with a low-rank decomposition:
$W+\Delta W=W+\Delta\alpha W_{\operatorname{down}}W_{\operatorname{up}},$ (7)
@@
$\theta^{*}=\theta+\Delta\theta,$ (8)
$\Delta\theta=\arg\min_{\Delta\theta}\sum_{(r,c^{*})}l_{ce}(\mathcal{M}_{\theta+\Delta\theta}(r),c^{*}).$ (9)
III-D Iterative Adaptation
The preceding three steps can go through multiple iterations until a certain number of rounds is reached or the performance no longer improves.
For $l$-th iteration that $l>1$, we initialize its initial model $\mathcal{M}_{\theta_{l}}$ as the enhanced model of the previous iteration $\mathcal{M}_{\theta^{*}_{l-1}}$. Based on $\mathcal{M}_{\theta_{l}}$, we repeat the process in steps of error code collection and automatic code revision to sample error codes $\{c^{\prime}\}_{l}$ and revised codes $\{c^{*}\}_{l}$, respectively. We use the union of $\{(r,c^{*})\}_{1:l}$, that is,
$\{(r,c^{*})\}_{1}\cup\cdots\cup\{(r,c^{*})\}_{i}\cdots\cup\{(r,c^{*})\}_{l},$ (10)
@@ III-D Iterative Adaptation
This iteration is a step-by-step optimization designed to continuously improve the adaptability of models to the specific scenario. The complete process of SEED is summarized in Algorithm 1.
Algorithm 1 Pseudocode of the proposed SEED.
0:  Few-sample Dataset $\mathcal{D}_{train}=\{(r,c)\}$, initial LLM $\mathcal{M}_{\theta}$.
@@ IV Evaluation
We present extensive experiments that span three representative code generation datasets, two fine-tuning settings, and five different LLMs of varying sizes or series. We aim to investigate six research questions:
  • RQ1: How does SEED perform compared to the mainstream adaptation approaches?
  • RQ2: How does SEED perform when applied to various LLMs?
  • RQ3: What kind of training sample has the better training effect?
  • RQ4: How does the number of iterations affect the effectiveness of SEED?
@@
  • RQ5: What is the impact of implementing the automatic code revision component of SEED in conjunction with alternative LLMs?
  • RQ6: How does each input component of Self-Revise contribute to the effectiveness?
IV-A Datasets
MBPP [26] contains crowd-sourced Python programming problems, covering programming fundamentals. We selected the version in the work [27], which consists of 276 problems and some generated error codes alongside their human-revised counterparts, thus facilitating subsequent experiments.
@@ IV-A Datasets
HumanEval [1] is a widely-used code generation benchmark, containing 164 handwritten programming problems, proposed by OpenAI. Each programming problem includes a function signature, a natural language description, use cases, a correct solution in Python, and several test tests.
@@
DS-pandas [28] comprises 291 data science problems utilizing Pandas libraries, sourced from real-world problems posted by developers on StackOverflow. This dataset can evaluate the ability of LLMs to utilize specific data-analysis libraries for code generation.
Due to the scarcity and privacy of domain data, it becomes difficult to obtain these data, and there are no available datasets. Therefore, we use the MBPP, HumanEval, and DS-pandas to simulate the specific scenario with limited samples. We choose the first 100 problems from all datasets as test dataset $\mathcal{D}{\text{test}}$. The remaining problems (up to 200 problems) serve as training dataset $\mathcal{D}{\text{train}}$. The statistics of all datasets are presented in Table I.
TABLE I: Statistics of MBPP, HumanEval, and DS-pandas.
Dataset Num of Problems Avg. Length Avg. Num
$\mathcal{D}_{\text{train}}$ $\mathcal{D}_{\text{test}}$ Requirement Code Test Case
MBPP 176 100 15.7 32.5 3
HumanEval 64 100 23.0 30.5 8
DS-pandas 191 100 184.8 26.2 2
IV-B Implementation Details
@@
We use a single A6000 GPU to conduct all experiments. We select CodeGen-2B [5] as our base model by default, which is a well-known open-source LLM for code and is suitable for full fine-tuning within our computational resource constraints. $\mathcal{M}_{\theta}$ is initialized to our base model, and $\mathcal{M}_{\operatorname{Revise}}$ is derived from $\mathcal{M}_{\theta}$ though Self-ReviseIII-B).
@@
For full parameter fine-tuning, i.e., Fine-tuning (Full) [24], we use the AdamW optimizer [29], with hyperparameters $\beta_{1}=0.9$ and $\beta_{2}=0.9$, accompanied by a linear learning rate schedule. The initial learning rate is set to 5e-6, with a batch size of 1 and gradient accumulation of 32 steps for training across 10 epochs. For parameter-efficient fine-tuning, i.e., Fine-tuning (LoRA) [25], the learning rate is set to 2e-4. Additionally, the rank $r$ is adjusted to 128, and the scaling factor $\alpha$ is set at 8. All other hyperparameters remain aligned with Fine-tuning (Full). For few-shot prompting [30], we set the number of examples in prompt to 4. All baselines in the experiments use consistent settings.
@@
In the error code collection step (§III-A) and the automatic code revision step (§III-B), we use temperature [31, 32] sampling to generate multiple samples: 5 samples in the former and 30 in the latter, with the temperature set to 0.8. To obtain the final revised code in the automatic code revision step, we choose the one of revised code exhibiting the minimum Levenshtein distance [33] to the error code. The number of iterative adaptations is set to 2. The maximum generation length is uniformly limited to 256 tokens.
IV-C Metrics
@@
Following the practice of real software development which utilizes testing for evaluation [34, 35], we employ the Pass@k [36] metric to measure the functional correctness of the generated code by executing test cases. We use the unbiased version [1] of Pass@k, where $n>=k$ samples are generated for each problem, count the number of correct samples $c<=n$ which pass test cases and calculate the following estimator,
@@
$\footnotesize{\operatorname{Pass@k}=\mathop{\mathbb{E}}\limits_{\operatorname{Problems}}\begin{bmatrix}1-\frac{\begin{pmatrix}n-c\\ k\end{pmatrix}}{\begin{pmatrix}n\\ k\end{pmatrix}}\end{bmatrix}.}$ (11)
@@
For automatic code revision, we add the pass@any metric which refers to the percentage of tasks for which the model generates at least one correct code that passed all test cases.
@@
In the final evaluation of this paper, we set the temperature to 0.8 and generate $n=50$ samples, which are then used to calculate unbiased Pass@k [1] via Eq. (11) in all experiments. All evaluation results are averaged over five test runs.
IV-D RQ1: Comparison of SEED and mainstream adaptation approaches
@@
Baselines. In this section, we evaluate the effectiveness of SEED by comparing it against other four mainstream approaches used as baselines:
@@
  • Fine-tuning (LoRA): We fine-tune the LLM on $\mathcal{D}_{train}$ using LoRA [25], which greatly reduces the number of trainable parameters.
  • Few-shot Prompting: We use 4-shot prompt [30] to align LLMs with the input-output format of $\mathcal{D}_{train}$, where 4 examples in prompt are randomly selected from $\mathcal{D}_{train}$.
Considering the baselines involve full-parameter fine-tuning, CodeGen-2B is uniformly selected as the base model in this experiment. For SEED, we use 30% data of $\mathcal{D}_{train}$ for Self-Revise (FT) [^4] In addition to MBPP dataset, for other two datasets (i.e., HumanEval and DS-pandas), we generate one error code per sample in a subset comprising 30% of the training set, using CodeGen-2B. Subsequently, authors collaboratively apply the minimal necessary revisions to correct these error codes., while the remaining 70% data of $\mathcal{D}_{train}$ is employed for model optimizing, where we use full-parameter fine-tuning.
TABLE II: Pass@k (%) of SEED and baselines on MBPP, HumanEval, and DS-pandas benchmarks, and the teal number after $\uparrow$ denotes the relative improvement of SEED over Fine-tuning (Full).
Benchmark Approach Pass@1 Pass@5 Pass@10
MBPP Direct Generation 15.6% 31.4% 40.2%
Fine-tuning (Full) 25.8% 46.2% 57.6%
Fine-tuning (LoRA) 19.8% 39.8% 55.2%
Few-shot Prompting 24.4% 38.0% 49.4%
[…]
@@
[…]
DS-pandas Direct Generation 0.8% 3.1% 5.6%
@@
[…]
@@
Results. We conducted experiments on three public benchmarks, i.e., MBPP, HumanEval, and DS-pandas. The experimental results are summarized in Table IV-D. This comparison yielded three insightful observations: 1) Significant superiority of SEED: Our proposed SEED performs significantly better than the other four baselines on the three benchmarks. Notably, SEED exhibits significant relative improvements of 27.2%, 32.9%, and 103.9%, respectively, when compared to the best-performing baseline Fine-tuning (Full). 2) Worst performance of Direct Generation: The performance of Direct Generation is significantly lower than the Fine-tuning (Full), Fine-tuning (LoRA), and Prompt baselines. This result suggests that directly applying LLMs for evaluation may be less suitable for specific scenarios, resulting in performance differences. 3) Fine-tuning (LoRA) is less effective than Fine-tuning (Full): Although Fine-tuning (LoRA) offers the advantage of reduced computational resource requirements for fine-tuning LLMs, it trades off the performance. 4) Less improvement of Few-shot Prompting: Few-shot Prompting generally underperforms compared to the Fine-tuning approaches. It only outperforms Fine-tuning (LoRA) in the MBPP benchmark’s Pass@1 metric. However, in other situations, the efficacy of Few-shot Prompting is the lowest, with the exception of Direct Generation.

### v3

@@ 3.3. Model Optimization
For fine-tuning (Devlin et al., 2019), we update all parameter $\theta$ of LLMs as
(7) $\theta^{*}=\arg\min_{\theta}\sum_{(r,c^{*})}l_{ce}(\mathcal{M}_{\theta}(r),c^{*}),$
When the computational resources are insufficient, we employ Low-Rank Adaptation (LoRA) (Hu et al., 2022) to fine-tune LLMs. For a weight matrix $W\in\mathbb{R}^{d\times k}$, LoRA represents its update with a low-rank decomposition:
(8) $W+\Delta W=W+\Delta\alpha W_{\operatorname{down}}W_{\operatorname{up}},$
@@
(9) $\theta^{*}=\theta+\Delta\theta,$
(10) $\Delta\theta=\arg\min_{\Delta\theta}\sum_{(r,c^{*})}l_{ce}(\mathcal{M}_{\theta+\Delta\theta}(r),c^{*}).$
3.4. Iterative Adaptation
The preceding three steps can go through multiple iterations until a certain number of rounds is reached or the revised code no longer increases.
For $l$-th iteration that $l>1$, we initialize its initial model $\mathcal{M}_{\theta_{l}}$ as the enhanced model of the previous iteration $\mathcal{M}_{\theta^{*}_{l-1}}$. Based on $\mathcal{M}_{\theta_{l}}$, we repeat the process in steps of error code collection and automatic code revision to sample error codes $\{c^{\prime}\}_{l}$ and revised codes $\{c^{*}\}_{l}$, respectively. Inspired by experience replay (Mnih et al., 2015) in reinforcement learning, we use the union of collected data in each previous iteration $\{(r,c^{*})\}_{1:l}$ to stabilize the learning process and improve data utilization efficiency, i.e.,
(11) $\{(r,c^{*})\}_{1}\cup\cdots\cup\{(r,c^{*})\}_{i}\cdots\cup\{(r,c^{*})\}_{l},$
@@ 3.4. Iterative Adaptation
This iteration is a step-by-step optimization designed to continuously improve the adaptability of models to the specific scenario. The complete process of DEED is summarized in Algorithm 1.
Algorithm 1 Pseudocode of DEED.
0:  Dataset $\mathcal{D}_{train}=\{(r,c)\}$, initial LLM $\mathcal{M}_{\theta}$.
@@ 4. Evaluation Setup
We present extensive experiments that span five representative code generation datasets, two fine-tuning settings, and four different LLMs of varying sizes or series. We aim to investigate six research questions:
  • RQ1: How does DEED perform compared to the mainstream adaptation approaches? This question aims to investigate the superiority of our method compared to existing approaches in data-scarce scenarios.
  • RQ2: How does DEED perform when applied to various LLMs? This question explores the usefulness of DEED across different LLMs, assessing its applicability and performance consistency in varying LLMs.
  • RQ3: What kind of training sample has the better training effect? This question helps us understand the promotion of training efficiency by training with the revision of LLMs’ erroneous outputs.
  • RQ4: How does the number of iterations affect the effectiveness of DEED? This question explores the changes in LLMs’ performance during the process of iterative optimization, to demonstrate the necessity of iterative adaptation in our method.
@@
  • RQ5: What is the impact of implementing the automatic code revision component of DEED in conjunction with alternative LLMs? We choose to use LLMs themselves for revision, termed Self-Revise. This question explores the effects of using other LLMs for revision.
  • RQ6: How does each input component of Self-Revise contribute to the effectiveness? Since Self-Revise utilizes components such as correct solutions, failed test cases, and error messages, this question aims to analyze the impact of these components on its performance.
4.1. Datasets
We use five public code generation datasets to simulate the specific scenario with limited data and apply DEED to each dataset to evaluate its effectiveness. Specifically,
@@ 4.1. Datasets
  • HumanEval (Chen et al., 2021) is a widely-used code generation benchmark, containing 164 handwritten programming problems, proposed by OpenAI. Each programming problem includes a function signature, a natural language description, use cases, a correct solution in Python, and several test cases.
  • MBPP (Austin et al., 2021) contains crowd-sourced Python programming problems. We use the version in the work (Chen et al., 2023c), which consists of 276 problems and some generated error codes alongside their human-revised counterparts, thus facilitating subsequent experiments.
  • HumanEval-ET and MBPP-ET (Dong et al., 2023a) are extended versions of MBPP and HumanEval, respectively, where each includes over 100 additional test cases per problem. This updated version enhances the soundness of code evaluation compared to the original benchmarks.
@@
  • DataScience (Lai et al., 2023) comprises 291 data science problems utilizing Pandas libraries. This dataset can evaluate the ability of LLMs to utilize specific data-analysis libraries for code generation.
We sample $\min(200,40\%*\mathcal{D})$ problems from the datasets as $\mathcal{D}_{{train}}$, while the remaining problems serve as $\mathcal{D}_{{test}}$.
4.2. Implementation Details
@@
We use a single A6000 GPU to conduct all experiments. We select CodeGen-2B (Nijkamp et al., 2023) as our base model by default, which is a well-known open-source LLM for code and is suitable for full fine-tuning within our computational resource constraints. $\mathcal{M}_{\theta}$ is initialized to our base model, and $\mathcal{M}_{\operatorname{Revise}}$ is derived from $\mathcal{M}_{\theta}$ though Self-Revise (§3.2). In fine-tuning setting, $\mathcal{M}_{\operatorname{Revise}}$ only needs to be trained at the beginning and then remains unchanged for subsequent operations.
@@
For full parameter fine-tuning, i.e., Fine-tuning (Full) (Devlin et al., 2019), we use the AdamW optimizer (Loshchilov and Hutter, 2017), with hyperparameters $\beta_{1}=0.9$ and $\beta_{2}=0.9$, accompanied by a linear learning rate schedule. The initial learning rate is set to 5e-6, with a batch size of 1 and gradient accumulation of 32 steps for training across 10 epochs. For parameter-efficient fine-tuning, i.e., Fine-tuning (LoRA) (Hu et al., 2022), the learning rate is set to 2e-4. Additionally, the rank $r$ is adjusted to 128, and the scaling factor $\alpha$ is set at 8. All other hyperparameters remain aligned with Fine-tuning (Full). For few-shot prompting (Brown et al., 2020a), we set the number of examples in prompt to 4. All baselines in the experiments use consistent settings.
@@
In the error code collection step (§3.1) and the automatic code revision step (§3.2), we use temperature (Holtzman et al., 2020; Ackley et al., 1985) sampling to generate multiple samples: 5 samples in the former and 30 in the latter, with the temperature set to 0.8. To obtain the final revised code in the automatic code revision step, we choose one of the revised codes exhibiting the minimum Levenshtein distance (Levenshtein et al., 1966) to the error code. The number of iterative adaptations is set to 2. The maximum generation length is uniformly limited to 1024 tokens.
4.3. Metrics
@@
Following the practice of real software development which utilizes testing for evaluation (Ruparelia, 2010; Abrahamsson et al., 2002), we employ the Pass@k (Li et al., 2022) metric to measure the functional correctness of the generated code by executing test cases. We use the unbiased version (Chen et al., 2021) of Pass@k, where $n>=k$ samples are generated for each problem, count the number of correct samples $c<=n$ which pass test cases and calculate the following estimator,
@@
(12) $\footnotesize{\operatorname{Pass@k}=\mathop{\mathbb{E}}\limits_{\operatorname{Problems}}\begin{bmatrix}1-\frac{\begin{pmatrix}n-c\\ k\end{pmatrix}}{\begin{pmatrix}n\\ k\end{pmatrix}}\end{bmatrix}.}$
@@
For automatic code revision, we add the pass@any metric which refers to the percentage of tasks for which the model generates at least one correct code that passes all test cases.
@@
In the final evaluation of this paper, we set the temperature to 0.8 and generate $n=50$ samples, which are then used to calculate unbiased Pass@k (Chen et al., 2021) via Eq. (12) in all experiments. All evaluation results are averaged over five test runs.
5. Results
5.1. RQ1: Comparison of DEED and the mainstream adaptation approaches
@@
Baselines. In this section, we evaluate the effectiveness of DEED by comparing it against four mainstream approaches used as baselines:
@@
  • Fine-tuning (LoRA): We fine-tune the LLM on $\mathcal{D}_{train}$ using LoRA (Hu et al., 2022), which greatly reduces the number of trainable parameters.
  • Few-shot Prompting: We use 4-shot prompt (Brown et al., 2020a) to align LLMs with the input-output format of $\mathcal{D}_{train}$, where 4 examples in prompt are randomly selected from $\mathcal{D}_{train}$.
  • Self-refine (Madaan et al., 2023) and Self-debug (Chen et al., 2023a) iteratively refine the generated code through prompting techniques. The number of iterations is set to 2.
Considering the baselines involve full-parameter fine-tuning, CodeGen-2B is uniformly selected as the base model in this experiment. For DEED, we use 30% data of $\mathcal{D}_{train}$ for Self-Revise (FT) [^1] In addition to MBPP dataset, for the other two datasets (i.e., HumanEval and DataScience), we generate one error code per sample in a subset comprising 30% of the training set, using CodeGen-2B. Subsequently, authors collaboratively apply the minimal necessary revisions to correct these error codes., while the remaining 70% data of $\mathcal{D}_{train}$ is employed for model optimizing, where we use full-parameter fine-tuning.
Table 1. Pass@k (%) of DEED and baselines on HumanEval, MBPP, and DataScience datasets. The values in brackets represent results on the extended version of the datasets, and the teal number after $\uparrow$ denotes the relative improvement of DEED over the second-highest score.
Datasets Method Pass@1 Pass@5 Pass@10
@@
Self-Refine 25.3% 45.2% 51.9%
Self-Debug 26.4% 46.4% 54.2%
DEED $\textbf{38.6\%}\ \ ({\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\uparrow 29.5\%})$ 54.7% 62.2%
HumanEval-ET Direct Generation 17.0% 30.8% 36.3%
Fine-tuning (Full) 21.5% 34.2% 41.5%
Fine-tuning (LoRA) 20.3% 33.7% 39.9%
Few-shot Prompting 19.1% 33.5% 39.1%
Self-Refine 17.5% 32.6% 38.2%
Self-Debug 19.2% 33.6% 41.0%
DEED $\textbf{28.6\%}\ \ ({\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\uparrow 33.0\%})$ 39.3% 46.5%
MBPP Direct Generation 15.6% 31.4% 40.2%
Fine-tuning (Full) 25.8% 45.2% 57.6%
Fine-tuning (LoRA) 19.8% 39.8% 55.2%
Few-shot Prompting 24.4% 38.0% 49.4%
Self-Refine 25.6% 38.8% 50.2%
Self-Debug 20.2% 34.5% 40.6%
DEED $\textbf{32.8\%}\ \ ({\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\uparrow 27.1\%})$ 46.8% 64.0%
MBPP-ET Direct Generation 10.6% 21.7% 28.2%
Fine-tuning (Full) 18.1% 33.1% 43.7%
Fine-tuning (LoRA) 15.1% 32.5% 42.4%
Few-shot Prompting 17.5% 27.3% 36.9%
Self-Refine 17.8% 27.6% 37.2%
Self-Debug 14.1% 25.7% 29.6%
DEED $\textbf{24.9\%}\ \ ({\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\uparrow 37.6\%})$ 36.3% 49.3%
DataScience Direct Generation 0.8% 3.1% 5.6%
@@
Self-Refine 2.1% 4.7% 5.8%
Self-Debug 1.2% 2.8% 4.1%
DEED $\textbf{5.3\%}\ \ ({\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\uparrow 103.8\%})$ 9.5% 12.3%
@@
Results. We conducted experiments on five public datasets, i.e., MBPP, HumanEval, and DataScience. The experimental results are summarized in Table 1. This comparison yielded four insightful observations: 1) Significant superiority of DEED: Our proposed DEED performs significantly better than the other six baselines on the five datasets. Notably, DEED exhibits significant relative improvements of 29.5%, 33.0%, 27.1%, 37.6%, and 103.8%, respectively, when compared to the best-performing baseline Fine-tuning (Full). Self-refine and Self-debug underperform on small LLMs like codegen-2B. Self-debug excels over Self-refine only on the HumanEval dataset, where public test cases and results are available. On other datasets, Self-debug relies on the LLM’s generated code explanations and feedback. Moreover, we find that DEED not only surpasses Self-refine and Self-debug in terms of performance but also in cost. These methods have a significant disadvantage in cost as they require iterative refinements for each sample, leading to increased time and token consumption. Their cost is directly proportional to the number of sampling and the number of iterations, while DEED is free of these two factors during inference. 2) Worst performance of Direct Generation: The performance of Direct Generation is significantly lower than Fine-tuning (Full), Fine-tuning (LoRA), and Prompt baselines. This result suggests that directly applying LLMs for evaluation may be less suitable for specific scenarios, resulting in performance differences. 3) Fine-tuning (LoRA) is less effective than Fine-tuning (Full): Although Fine-tuning (LoRA) offers the advantage of reduced computational resource requirements for fine-tuning LLMs, it trades off the performance. 4) Less improvement of Few-shot Prompting: Few-shot prompting is the most commonly used prompting technique, but its main limitation lies in its difficulty in imparting new knowledge or developing new capabilities in the model. It primarily assists the model in adjusting its outputs to better align with expected results, therefore its adaptability is limited.
{answerbox}